OpenStudy (anonymous):

Given the function f(x)=5x-9, find the set of values of x such that f(x) = 6. Justify your solution by using the properties of equality

OpenStudy (mathmale):

Too bad you chose that user name, and too bad that OpenStudy won't let you change it. But anyway. Simply replace f(x) with 6 and equate that to 5x-9. Solve the resulting equation.

OpenStudy (mathmale):

Hint: add 9 to both sides, and then simplify the equation.

OpenStudy (anonymous):

f(6)=5(6)-9 f(6)=30-9=21 the f(x) just needs you to change it :) nicce name btw lol

OpenStudy (mathmale):

MathWhiz (your new name): Let's start over. f(x)=5x-9 But we are told that f(x)=6. f(x) here is a label. Then replace f(x) with 6. 6=5x-9 Add 9 to both sides. You do it, please.

OpenStudy (the_fizicx99):

@hayleya0202 Your answer is incorrect.

OpenStudy (mathmale):

We are told that f(x)=6. We are not told to let x=6 within f(x).

OpenStudy (the_fizicx99):

Yes, as my bff said ^